what volume of oxygen is required to produce 300cm3 of SO3 at room temperature and pressure
1
Expert's answer
2020-07-24T14:52:31-0400

Balance the equation

2S+3O2=2SO3


300 ml is the same as 0.3 litres

thus the moles per volume and ideal pressure will be

n(SO3)=V(SO3)/Vm=0,3/22,4=0.013mol

multiply by 3 due to volume ratio to get the number of moles of oxygen.


n(O2)=0.013*3/2=0.0195mol

multiply by ideal pressure volume of gases at room temperature to get the volume of oxygen required.

V(O2)=0.0195×22,4=0.437l=437ml


therefore volume is approximately 437ml